From GlueXWiki
Jump to: navigation, search

October 28, 2010 FDC meeting

Tentative Agenda

  1. Production (David)
  2. Electronics (Fernando, Chris, Roger)
    • Update on PCBs
    • Cathode foil and rigid-flex procurement
    • fADC125/F1TDC synchronization (Beni, Gerard) FDC Log Book, page 531
  3. Engineering (Bill, David)
    • Wire stringing
    • Cathode foils: plans for inspection, storage, and cathode lamination
  4. Full-scale prototype tests (Lubomir)
  5. Other


Participants: Fernando, Eugene, Casey, Bill, Mark, David, Chris, Simon, Glenn, Roger, Beni, Gerard (on the phone) and Lubomir.


- The humidity was very high yesterday (above 80%) at Blue Crab and the Rohacell material that was not laminated deformed (photos at page 533). We still have to check if the four laminated wire frames that are at Blue Crab are OK. Bill: we can't stop the production and should continue with the lamination. Slution: to put a dehumidifier at one of the offices at Blue Crab and to use it for the Rohacell lamination. To recover the original Rohacell shape ones needs to heat it at 250degF, but don't have to do it because we have enough material (Casey).

- We will try to do both the Rohacell ring gluing and the lamination to the wire frame at once with vacuum bagging. Bill is thinking about a box for storage flushed with nitrogen. 14 wire frames (almost half) are laminated.


- Chris was investigating different ways of cleaning the PCBs using Vigor flux-off. Bill asked if we can buy an ultrasonic cleaner 42"x18"x6" that will be used also to clean the wire frames after soldering the wires, for $5500. Glenn agreed.

- Fernando: the first article signal cables will arrive by mid November and Chris is working on a board to be used for testing. The solder material (japanese) recommended from UVA is no longer available. Chris ordered similar materialial from other company to exercise. Working on the sole source procurement of the ASICs. Fermilab is also intersted in the same ASICs and Fernando discussed if we can share with Fermilab the costs for the mask (~$100K).

- The contract for the cathode foils was awarded. Alflex is waiting for the Gerber files from Roger; the files are ready waiting final approval from Bill and Fernando. After that we will have the first article in three weeks, then we have one week (around Thanksgiving) to approve it before the start of the whole production.

- Rigid-flex: Roger is waiting for quotations from two companies after some additional explanation concerning the coating. Allflex already decided not to bid for this. Then Roger will submit a PR for 500 rigid-flexes. Once more we discussed the technique to be used for soldering to the cathode foils. Fernando prefers to use conventional soldering instead of using conductive film; the argument: the later technology exists for only 4 years and we are not sure if the soldering will work for 10 years. We will have plenty of material to exercise this and to measure the failure rate of the standard method and to try the new one. After that we should make a decision.

- Beni showed a drift time distribution (page 531) as measured with fADC125. It has 8ns structure coming form the 125MHz clock. Gerard doubts that such structure should appear. Beni, Gerard, and Fernando discussed the ways one can measure the time between the trigger and the phase of the fADC clock that is needed to improve the timing resolution from the fADC. In the experiment the clock will be common and a special module will provide this opportunity. For now, Gerard suggested plugging the trigger into the fADC to measure the trigger time.


- Wire stringing: David will try tomorrow to use the tension laser system to find the wire position. Al made progress in the USB connection to the stepper motor: now USB can talk to the ISA card. Bill accepted different proposals how the make the weights for the wire tensioning.

- We started discussing the acceptance procedure of the cathode foils. The first article will arrive around Thanksgiving and the problem is to find clean space at that time to do the inspection and some tests. A possible solution will be to use some space in the Hall B clean room (EEL building). Lubomir will talk to Steve Christo and then to Mac Mestayer. Eugene: chances are not very high and Roger will postpone submitting the Gerber files till we have a solution.

Full-scale prototype tests

- The gas gain was measured using two different methods and compared to the Magboltz (inside Garfield) calculations (page 530). Both methods use 55Fe source. In the first, one measures the rates on the wires and relates it to the total current. In the second, one uses Garfield to simulate the signals, including the measured electronics response, except for the gain that is adjusted to the measured amplitudes. The two measurements are consistent but different from the Magboltz calculations, by up to a factor of 10 for 90/10.

- The old problem (page 499) of the ratio top/bottom cathode signal being > 1 was re-visited (page 531) now for different gas mixture/high voltages. The effect increases by increasing the HV, suggesting electrostatic forces to be responsible. Other possibilities were discussed; no explanation yet.